# Intuitive Fine-Tuning: Towards Simplifying Alignment into a Single Process

### v2

@@ 5.1 Settings for NLP
Datasets. We select UltraChat-200k ding2023enhancing and UltraFeedback-60k cui2023ultrafeedback as single-target and pair-wise dataset respectively.
Models. We conduct experiments on Mistral-7B-v0.1 jiang2023mistral and Mistral-7B-sft-beta tunstall2023zephyr , with the former as the base model and the latter fine-tuned on UltraChat-200k.
@@
Baselines. Our main baselines are SFT and DPO rafailov2024direct , with PPO excluded due to computational limitations. Additionally, we incorporate an improved versions of DPO, named ORPO hong2024reference , who claims to achieve alignment directly without SFT and reference model, aligning well with our goals. In addition to reproducing the algorithms mentioned above, we also consider Zephyr-7B-beta tunstall2023zephyr and Mistral-ORPO-alpha hong2024reference , two open-source checkpoints that utilize sequential and direct recipes respectively. Both of them used start-point models and datasets similar to ours.
Benchmarks. We consider two types of benchmarks. One is from the widely used Open-LLM LeaderBoard, which contains ARC-Challenge (25-shot) clark2018think , MMLU (5-shot) chung2024scaling , TruthfulQA (0-shot) lin2021truthfulqa , WinoGrande (5-shot) sakaguchi2021winogrande , and GSM8K (5-shot) cobbe2021training . The other is LM-based evaluation, including Alpaca-Eval and Alpaca-Eval-2 dubois2024alpacafarm . We utilize chat template for all benchmarks to obtain a more accurate evaluation for chat models.
@@ 5.2 Main Results in NLP Tasks
Effectiveness on Sequential Recipe. In this scenario, IFT demonstrates good performance across benchmarks having standard answers or not, see more details in Table 2 and 3. On Open-LLM Leaderboard, IFT showcases the best average capabilities across all tasks, excelling particularly in tasks requiring generation, reasoning and fact-following abilities, such as TruthfulQA and GSM8K. However, IFT has a relatively large gap between DPO in multi-choice tasks like ARC-Challenge and MMLU. We will explained the reasons later. When evaluated for instruction-following and question-answering abilities as judged by GPT-4, IFT’s performance is comparable to that of the chosen baselines. Remarkably, IFT achieves these results using the least amount of data and computational resources among all the methods tested.
Table 2: Evaluation on LLM-based Benchmarks. In the typical sequential setting, SFT+DPO still achieves the highest scores, although IFT performs comparably. However, when aligning directly from the base model, IFT performs not only superiorly to other methods but also on par with sequential recipes. As for open-source models, we directly present its publicly available score.
Method Reference Data Alpaca-Eval Alpaca-Eval-2 pairwise volume win-rate lc win-rate win-rate lc win-rate Mistral-7B 24.72 11.57 1.25 0.35      fine-tuning with UltraFeedback-60k + SFT 120k 82.56 78.32 7.09 8.67 + DPO 120k 74.00 73.12 9.73 8.58 + ORPO 120k 85.14 76.60 8.82 12.34 + IFT 120k 85.18 78.78 9.95 13.27 Mistral-ORPO-$\alpha$ 120k 87.92 11.33      fine-tuning with UltraChat-200k + UltraFeedback-60k sequentially + SFT 200k 86.69 77.96 4.08 6.43 + SFT + SFT 260k 86.34 76.98 4.55 7.14 + SFT + DPO 320k 91.62 81.54 10.08 13.72 + SFT + ORPO 320k 86.26 79.67 7.40 12.27 + SFT + IFT 260k 88.37 81.29 10.26 14.34 Zephyr-7B-$\beta$ 320k 90.60 10.99
Method ARC ARC-Gen MMLU TruthfulQA WinoGrande GSM8K Avg. Mistral-7B 53.07 73.04 59.14 45.29 77.58 38.89 54.79      fine-tuning with UltraFeedback-60k + SFT 56.49 74.00 60.44 55.57 77.90 42.84 58.65 + DPO 61.86 73.54 61.02 47.98 76.64 43.89 58.28 + ORPO 56.66 73.98 60.57 51.77 77.19 42.30 57.70 + IFT 56.74 74.15 60.49 57.65 78.45 44.73 59.61 Mistral-ORPO-$\alpha$ 57.25 73.72 58.74 60.59 73.72 46.78 59.41      fine-tuning with Ultrachat-200k + UltraFeedback-60k sequentially + SFT 57.68 72.87 58.25 45.78 77.19 40.94 55.97 + SFT + SFT 58.10 72.61 58.40 48.59 76.80 43.06 56.99 + SFT + DPO 63.91 73.98 59.75 46.39 76.06 41.47 57.52 + SFT + ORPO 58.45 73.21 58.80 50.31 76.45 42.76 57.35 + SFT + IFT 58.36 73.38 58.45 52.39 78.06 43.82 58.22 Zephyr-7B-$\beta$ 67.41 72.61 58.74 53.37 74.11 33.89 57.50
@@
Effectiveness of Preference Optimization Alone. IFT not only maintains the performance advantages compared with other baselines in this setting, as seen in the sequential scenario. But also, IFT performs comparably or even superiorly to many method in sequential recipe. While DPO tends to fail under this setting, ORPO remains competitive in its open-source model. However, when constrained in the same experiment setting, the performance of ORPO becomes worse than IFT. Additionally, the reliance on preference data makes ORPO more costly in terms of negative sampling, preference labeling, and GPU memory consumption. Consequently, IFT stands out as a more efficient and cost-effective alternative in this context.
@@
Multi-Choice vs. Generation. IFT performs better on generation tasks but struggles with multi-choice, whereas DPO exhibits the opposite performance. This may due to differences in evaluation metrics and training objectives zheng2023large ; plaut2024softmax ; tsvilodub2024predictions . Multi-choice tasks evaluate log-likelihood for entire answers, while generation tasks require token-by-token construction for causality and reasoning. DPO aligns the mapping between instructions and complete answers, while IFT emphasizes token-level causal relationships. As a result, DPO tends to excel in multi-choice tasks, while IFT performs better in token-by-token exploration tasks. In an ARC-Challenge adaptation to generation tasks, IFT demonstrates superiority without changing the benchmark’s distribution. Overall, IFT showcases its balanced performance across diverse tasks and achieving the highest average score.
@@
Objective Trade-off between SFT and Preference Optimization. Traditional Preference Optimization methods deliver excellent alignment performance, particularly in enhancing the instruction-following ability of language models, as showed in Table 2. However, fitting the different objectives of SFT and Preference Optimization involves trade-offs tunstall2023zephyr . Even slight overfitting on SFT may result in reduced effectiveness of Preference Optimization. This phenomenon is also observed in Table 3, where the models trained by sequential recipe of SFT and other Preference Optimization methods showcase obvious inferior results on Open-LLM Leaderboard even worse than SFT alone. Avoiding this trade-off, ORPO and IFT can achieve better and more stable performance by directly conducting alignment on the base model.
@@
Efficiency and Scaling Potential of IFT. Although IFT achieves comparable or superior performance to other methods, it also boasts high efficiency in many aspects. Like SFT and ORPO, IFT does not require a reference model, which conserves GPU memory and computational resources. Most importantly, IFT and SFT are the only methods that conduct alignment without preference data, offering significant benefits as follows. Firstly, this characteristic eliminates the need for synchronous storage and computation of pairwise data on the GPU, thereby reducing memory consumption and training duration. Secondly, negative sampling from models and human preference-labeling are no longer necessary, eliminating the highest cost associated with alignment, which has been a discarded but fundamental challenge in research so far. Furthermore, using only the target answer brings the potential for scaling in alignment process, mirroring the core benefits found in pre-training.
@@ 5.3 Further Validation in Frozen-Lake Environment
Figure 3: The Frozen Lake Game. Considering the MSE distance of parameters between the trained and optimal policy, IFT performs much better than SFT and ORPO, but slightly worse than DPO.
As scores on Open-LLM Leaderboard only partially reflect models’ performance, and GPT-4 inadequately models human language generation, further comparison to a truly optimal policy is necessary. Given the difficulty of obtaining an optimal policy representing human language, we validate our algorithm in a simplified setting called Frozen Lake frozenlake . In this environment, an agent attempts to find a gift on a nearly frozen lake with several holes, terminating the game upon finding the gift or falling into a hole. The limited number of states and actions in this game allows the optimal policy to be easily derived using classical RL methods.
To simulate parameterized policy alignment, we employ a two-layer fully connected neural network and design the environment with one optimal and one sub-optimal trajectory. The optimal parameterized policy is trained using the previously obtained optimal state-action probabilities, and various fine-tuning methods from LMs are compared. We evaluate performance by measuring the MSE distance between the optimal and trained policy parameters.
@@ 6 Conclusion
In this paper, we first interpret SFT and some typical Preference Optimization methods into a unified framework using Preference Estimation and Transition Optimization. We then introduce an efficient and effective method called Intuitive Fine-Tuning (IFT), which achieves alignment directly from the base model using non-preference-labeled data. Finally, experiments on widely used benchmarks in NLP setting and in Frozen Lake environment demonstrate the competitive performance of IFT.
Limitations and Future Work. Our validation of IFT is limited to the fine-tuning stage, where data volume is constrained, leaving the scalability of IFT unexplored. Additionally, we primarily use Mistral-7B for baseline testing, and the generalization of IFT to larger and more diverse models requires exploration in the future.
Acknowledgments and Disclosure of Funding
Special thanks to Yihao Liu, Che Jiang, and Xuekai Zhu for contributing to idea formulation, Pengfei Li, Xuanqi Dong, and Jingkun Yang for their assistance with mathematical definitions, Hong Liu, and Chushu Zhou for insightful discussions during the experiments.
@@ A.2 Loss Function of IFT
Let’s begin by considering only one sampled state $s_{n}^{*}$ constrained by $\rho_{0}$,
Given the assumption of $T^{*}(s_{n}^{*},\rho_{0})=T^{*}(s_{N}^{*},\rho_{0})=1$, we have:
$\mathcal{L}(\hat{T_{\theta}}(s_{n}^{*},\rho_{0}))$ $=\log\frac{\hat{T_{\theta}}(s_{n}^{*},\rho_{0})}{T^{*}(s_{n}^{*},\rho_{0})}$ (24)
$=\log\frac{\hat{T_{\theta}}(s_{n}^{*},\rho_{0})}{T^{*}(s_{N}^{*},\rho_{0})}$
$=\log\frac{1}{\hat{T_{\theta}}(s_{N}^{*},s_{n}^{*})}$
$=-\log\prod\limits_{i=n}^{N}\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))$
$=-\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))$
Next, we extend this to entire trajectories and the entire dataset to obtain the expected loss:
$\mathcal{L}_{\text{IFT}}(\mathcal{T}_{\theta},\delta_{\theta})$ $=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{n}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}\left[\sum\limits_{n=0}^{N}\mathcal{L}\big{(}\hat{T_{\theta}}(s_{n}^{*},\rho_{0})\big{)}\right]$ (25)
$=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}\left[-\sum\limits_{n=0}^{N}\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\right]$
where $s_{0}^{*}=\rho_{0}$, and $\mathcal{S}_{\rho_{0}}^{*}$ is the optimal $\rho_{0}$-constrained state space.

### v3

@@ 5.1 Settings for NLP
Datasets. Our main experiments use UltraChat-200k (Ding et al., 2023) as the single-target dataset and UltraFeedback-60k (Cui et al., 2023) as the pairwise preference dataset. We also include a variant of UltraFeedback-60k introduced by Meng et al. (2024), which is sampled from Gemma2 and LLaMA3 and labeled with preferences using ArmoRM.
Models. Our main experiments are conducted on Mistral-7B-v0.1 (Jiang et al., 2023) and Mistral-7B-sft-beta (Tunstall et al., 2023), with the latter one fine-tuned from the former using UltraChat-200k. We also consider models with different architectures and parameter scales, including Gemma-2B (Team et al., 2024) and LLaMA3-8B (Grattafiori et al., 2024).
@@
Baselines. SFT and DPO (Rafailov et al., 2024) are our main baselines, and we exclude PPO due to computational limitations. We also incorporate three improved versions of DPO: TDPO (Zeng et al., 2024), ORPO (Hong et al., 2024), and SimPO (Meng et al., 2024). TDPO transformers the DPO loss into token-level to make its objective closer to SFT. SimPO adds on a length-normalization term to replace the regularization from reference model. ORPO adds the SFT loss and a DPO-like loss together, achieving alignment directly without SFT and reference model. In addition to reproducing the algorithms mentioned above, we also consider Zephyr-7B-beta (Tunstall et al., 2023) and Mistral-ORPO-alpha (Hong et al., 2024), two open-source checkpoints that utilize sequential and direct recipes respectively. Both of them used start-point models and datasets similar to ours.
Benchmarks. We consider two types of benchmarks. One is from the widely used Open-LLM LeaderBoard, which contains ARC-Challenge(25-shot) (Clark et al., 2018), MMLU(5-shot) (Chung et al., 2024), TruthfulQA(0-shot) (Lin et al., 2021), WinoGrande(5-shot) (Sakaguchi et al., 2021), and GSM8K(5-shot) (Cobbe et al., 2021). The other is LM-based evaluation, including TL;DR (Völske et al., 2017), Alpaca-Eval, and Alpaca-Eval-2 (Dubois et al., 2024). As for TL;DR, we keep the same setting as (Rafailov et al., 2024), using GPT-4 to judge the win-rate between model’s generation and ground truth answer. We utilize chat template for all benchmarks to obtain a more accurate evaluation for chat models.
@@ 5.2 Main Results in NLP Tasks
Effectiveness on Sequential Recipe. In this scenario, IFT demonstrates good performance across benchmarks having standard answers or not (See Table 2 and 3 for details). On Open-LLM Leaderboard, IFT showcases the best average capabilities across all tasks, excelling particularly in tasks requiring generation, reasoning and fact-following abilities, such as TruthfulQA and GSM8K. However, IFT has a relatively large gap between DPO in multi-choice tasks like ARC-Challenge and MMLU. When evaluated for conversation and summarization judged by GPT-4, IFT’s performance is comparable to that of the chosen baselines. Remarkably, IFT achieves these results using the least amount of data and computational resources among all the methods tested.
@@
Effectiveness of Preference Optimization Alone. IFT not only maintains the performance advantages compared with other baselines in this setting. But also, IFT performs comparably or even superiorly to many method in sequential recipe (See Table 2, 3, and Appendix C for details). While DPO, SimPO and TDPO tend to fail under this setting, ORPO remains competitive in its open-source model. However, when constrained in the same experiment setting, the performance of ORPO becomes worse than IFT. Additionally, the reliance on preference data makes ORPO more costly in terms of negative sampling, preference labeling, and GPU memory consumption. Consequently, IFT stands out as a more efficient and cost-effective alternative in this context.
@@
Multi-Choice vs. Generation. IFT performs better on generation tasks but struggles with multi-choice, whereas DPO exhibits the opposite performance. This may due to differences in evaluation metrics and training objectives (Zheng et al., 2023; Plaut et al., 2024; Tsvilodub et al., 2024). Multi-choice tasks evaluate log-likelihood for entire answers, while generation tasks require token-by-token construction for causality and reasoning. DPO aligns the mapping between instructions and complete answers, while IFT emphasizes token-level causal relationships. As a result, DPO tends to excel in multi-choice tasks, while IFT performs better in token-by-token exploration tasks. In an ARC-Challenge adaptation to generation tasks, IFT demonstrates superiority without changing the benchmark’s distribution. Overall, IFT showcases its balanced performance across diverse tasks and achieving the highest average score.
@@
Objective Trade-off between SFT and Preference Optimization. Traditional Preference Optimization methods deliver excellent alignment performance, particularly in enhancing the instruction-following ability of language models, as showed in Table 3. However, fitting the different objectives of SFT and Preference Optimization involves trade-offs (Tunstall et al., 2023). Even slight overfitting on SFT may result in reduced effectiveness of Preference Optimization. This phenomenon is also observed in Table 2, where the models trained by sequential recipe of SFT and other Preference Optimization methods showcase obvious inferior results on Open-LLM Leaderboard even worse than SFT alone. Avoiding this trade-off, ORPO and IFT can achieve better and more stable performance by directly conducting alignment on the base model.
@@
Efficiency and Scaling Potential of IFT. Although IFT achieves comparable or superior performance to other methods, it also boasts high efficiency in many aspects. IFT does not require a reference model, which conserves GPU memory and computational resources. Most importantly, IFT and SFT are the only methods that conduct alignment without preference data, offering significant benefits as follows. Firstly, this characteristic eliminates the need for synchronous storage and computation of pairwise data on the GPU, thereby reducing memory consumption and training duration. Secondly, negative sampling from models and human preference-labeling are no longer necessary, eliminating the highest cost associated with alignment, which has been a discarded but fundamental challenge in research so far. Furthermore, using only the target answer brings the potential for scaling in alignment or even in pre-training.
@@ 5.3 Further Validation in Frozen-Lake Environment
Figure 3: The Frozen Lake Game. Considering the MSE distance between transition matrices of the trained and optimal policy, IFT performs much better than SFT and ORPO, but slightly worse than DPO.
As scores on Open-LLM Leaderboard only partially reflect models’ performance, and GPT-4 inadequately models human language generation, further comparison to a truly optimal policy is necessary. Given the difficulty of obtaining an optimal policy representing human language, we validate our algorithm in a simplified setting called Frozen Lake (Farama, 2023). In this environment, an agent attempts to find a gift on a nearly frozen lake with several holes, terminating the game upon finding the gift or falling into a hole. The limited number of states and actions in this game allows the optimal policy to be easily derived using classical RL methods.
To simulate parameterized policy alignment, we employ a two-layer fully connected neural network and design the environment with one optimal and one sub-optimal trajectory. The optimal parameterized policy is trained using the previously obtained optimal state-action transition matrix, and various fine-tuning methods from LMs are compared. We evaluate performance by measuring the MSE distance between the transition matrices of the optimal and trained policy. We didn’t count in TDPO and SimPO, as their objectives are similar as DPO in Frozen Lake Game.
@@
6 Related Work
Classical Reinforcement learning (RL) has demonstrated strong performance in various sequential decision-making and optimal control domains, including robotics (Levine et al., 2018), computer games (Vinyals et al., 2019) and others (Guan et al., 2021). There are two main categories of RL algorithms: value-based and policy-based, depending on whether they learn a parameterized policy. Value-based RL aims to fit an value function defined by Bellman Equation, containing methods such as Monte-Carlo (MC) Learning (Lazaric et al., 2007) and Temporal Difference Learning (Sutton, 1988; Seijen and Sutton, 2014). However, value-based methods struggle in continuous or large discrete space for its greedy objective. Thus, policy-based methods were introduced to model the decision-making process using a parameterized policy. As one of its best-known algorithms, Proximal Policy Optimization (PPO) (Schulman et al., 2017) is widely used in various domains, including Natural Language Processing (NLP).
Alignment for LMs has emerged as a crucial task these years, which adjusts the LMs’ generation distribution in line with human preferences (Bradley and Terry, 1952; Ziegler et al., 2019; Ouyang et al., 2022; Lee et al., 2023). While PPO remains the primary algorithm for alignment, its high demands for computation and memory hinders its broader use. Consequently, many improved methods have been proposed (Dong et al., 2023; Yuan et al., 2023; Zhao et al., 2023). Among them, DPO (Rafailov et al., 2024) unifies reward modeling and policy optimization by utilizing a loss function derived from PPO, training a single model to serve as both a policy model and a reward model. Without sacrificing performance, DPO decrease the costly consumption of PPO through directly value iteration similar to a preference-based format of MC instead of TD. However, it still relies on an expensive preference-labeling process and requires an SFT-based warm-up stage, which may introduce trade-offs when aligning the objectives of SFT and Preference Optimization.
Improved Versions of DPO come out one after another. Efforts such as (Liu et al., 2023b; Khaki et al., 2024; Yin et al., 2024; Guo et al., 2024b; Bansal et al., 2024; Liu et al., 2024) try to enhance the contrastive learning by utilizing better ranking strategies, more informative data, or more number of negative samples. Except for using offline data, (Liu et al., 2023a; Yuan et al., 2024; Guo et al., 2024a; Calandriello et al., 2024; Chen et al., 2024; Mitra et al., 2024) focus on online sampling and automated label/reward collection, reducing the manual cost required for alignment. Methods like (Ethayarajh et al., 2024; Hong et al., 2024) aim to reduce DPO’s dependency on SFT warm-up by transforming its loss functions and data format into a SFT manner. These algorithms handle positive and negative samples using SFT objective and Unlikelihood Training (Welleck et al., 2019), respectively. Recently, (Zeng et al., 2024; Meng et al., 2024) improved the integration of the SFT and DPO by introducing various regularization terms. These terms prevent the policy model from overfitting DPO objective and deviating from SFT objective. However, the actual volume of training data is not decreased in these methods. Also, GPU-memory-consuming pair-wise data is still required, while the need for a reference model and preference-labeling for the entire answer trajectory is only eliminated in limited cases.
@@ 7 Conclusion
In this paper, we first interpret SFT and typical Preference Optimization methods into a unified framework using Preference Estimation and Transition Optimization. Through this modeling, we found the biased prior used in SFT is one of the main reasons why SFT performs worse than other Preference Optimization methods. Then, we introduce an efficient and effective method called Intuitive Fine-Tuning (IFT), which achieves alignment directly from the base model using non-preference-labeled data. Finally, experiments on widely used NLP benchmarks and Frozen Lake environment demonstrate the competitive performance of IFT.
8 Limitations
Our validation of IFT is limited to the fine-tuning setting, where data volume is constrained, leaving the scalability of IFT unexplored.
Acknowledgments
This work is supported by the National Science and Technology Major Project (2023ZD0121403), and the Beijing Natural Science Foundation (IS23059). We further extend our gratitude to Yue Yu, Yihao Liu, Che Jiang, Xuekai Zhu, Jingkun Yang, Xuanqi Dong, Hong Liu, and Chushu Zhou for their insightful discussions with us.
@@ A.2 Loss Function of IFT
The disparities of transition between model and human can be formalized as follows:
$\mathcal{L}(\hat{T_{\theta}};T^{*})$ $=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{n}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}$ (23)
$\left[-\sum\limits_{n=0}^{N}\log\frac{\hat{T_{\theta}}(s_{n}^{*},\rho_{0})}{T^{*}(s_{n}^{*},\rho_{0})}\right]$
We make the same hypothesis as SFT that the optimization objective of each target intermediate state has a probability equal to 1, so that $\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}$:
$T^{*}(s_{n}^{*},\rho_{0})=1=T^{*}(s_{N}^{*},\rho_{0})$ (24)
Thus, the objective of IFT can be represented directly by the following loss function:
$\mathcal{L}(\hat{T_{\theta}})$ $=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}$ (25)
$\left[-\sum_{n=0}^{N}\log\mathcal{T}_{\theta}\left(\pi^{*}\left(\delta^{*}(s_{i}^{*})\right),\delta_{\theta}(s_{i}^{*})\right)\right]$
As the optimal policy enjoys the optimal transition:
$s_{i}^{*}=[s_{i-1}^{*},a_{i}^{*}]=[s_{i-1}^{*},\pi^{*}(s_{i-1}^{*})]=\Pi^{*}(s_{i-1}^{*})$ (26)
Therefore, the disturbed optimal state keeps similar with the original optimal state:
$\delta^{*}(s_{i}^{*})=(1-\lambda)s_{i}^{*}+\lambda\Pi^{*}(s_{i-1}^{*})=s_{i}^{*}$ (27)
Then, the final loss function can be presented as:
$\mathcal{L}(\mathcal{T}_{\theta},\delta_{\theta})$ $=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}$ (28)
$\left[-\sum\limits_{n=0}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\right]$